# STAR-RIS Enhanced UAV-Enabled MEC Networks with Bi-Directional Task Offloading

### v1

@@ II-B Task Offloading and Computation Model
where $p$ denotes the unified transmitted power of users, $B$ is the bandwidth of the system, $\sigma^{2}_{\mathrm{ua}}$ and $\sigma^{2}_{\mathrm{BS}}$ respectively denote the noise power at the BS and the UAV. In addition, $\boldsymbol{\Theta}_{\kappa}[n]=\operatorname{Diag}\big{\{}\beta^{1}_{\kappa}[n]e^{j\phi^{1}_{\kappa}[n]},\cdots,\beta^{m}_{\kappa}[n]e^{j\phi^{m}_{\kappa}[n]},\cdots,\beta^{M}_{\kappa}[n]e^{j\phi^{M}_{\kappa}[n]}\big{\}}$ is the matrix of the STAR-RIS’s coefficients with $\kappa\in\{\mathrm{r},\mathrm{t}\}$ representing the reflected or transmitted coefficients of the STAR-RIS, where the amplitudes $\beta^{m}_{\kappa}[n]$ and phases $\phi^{m}_{\kappa}[n]$ of the STAR-RIS should satisfy $(\beta^{m}_{\mathrm{r}}[n])^{2}+(\beta^{m}_{\mathrm{t}}[n])^{2}=1$, $\phi^{m}_{\mathrm{r}}[n],\phi^{m}_{\mathrm{t}}[n]\in[0,2\pi],\forall m\in\mathcal{M}\triangleq\{1,\cdots,m,\cdots,M\}$.
We assume that the ground users are with limited resources for local computing and thus users’ tasks have to be offloaded to the BS and UAV for computing. Let $l^{\mathrm{BS}}_{k}[n]$ and $l^{\mathrm{ua}}_{k}[n]$ denote the number of the offloaded bits that need to be computed at the BS and UAV for user $k$ in the $n$-th time slot, respectively. Due to the fact that the BS and UAV can only deal with the tasks they have received, the following constraints should be satisfied
$\delta_{\mathrm{t}}\zeta_{k}[n]R^{\mathrm{ua}}_{k}[n]\geq l^{\mathrm{ua}}_{k}[n],~{}\forall k\in\mathcal{K},n\in\mathcal{N},$ (14)
$\delta_{\mathrm{t}}\zeta_{k}[n]R^{\mathrm{BS}}_{k}[n]\geq l^{\mathrm{BS}}_{k}[n],~{}\forall k\in\mathcal{K},n\in\mathcal{N}.$ (15)
In order to ensure that every user’s minimum computational need is met, we implement the following QoS constraint
@@
where $L_{k}$ is the minimum computing tasks of the $k$-th user.
Let $f_{k}^{\mathrm{BS}}[n]$ and $f_{k}^{\mathrm{ua}}[n]$ respectively denote the computation frequency allocated by the BS and UAV for the $k$-th user at the $n$-th time slot, we have
@@
where $F_{\mathrm{BS}}$ and $F_{\mathrm{ua}}$ are the maximum CPU frequency at the BS and the UAV, respectively. In addition, considering the processing causality, it is assumed that the BS and the UAV solely receive the offloaded tasks without carrying out any computations within the first time slot, and the users cease the act of offloading their tasks at the last time slot, i.e., $f^{\mathrm{BS}}_{k}[1]=f^{\mathrm{ua}}_{k}[1]=l^{\mathrm{BS}}_{k}[N]=l^{\mathrm{ua}}_{k}[N]=0$ for $k\in\mathcal{K}$. In order to guarantee that all users’ offloaded task-input data can be completely computed within the give mission time $T$, we have the following information-causality constraints:
$\sum\limits_{i=2}^{n}\frac{f^{\mathrm{ua}}_{k}[i]\delta_{\mathrm{t}}}{\varrho_{\mathrm{ua}}}\geq\sum\limits_{i=1}^{n-1}l^{\mathrm{ua}}_{k}[i],~{}\forall k\in\mathcal{K},n\in\mathcal{N}_{1},$ (18)
$\sum\limits_{i=2}^{n}\frac{f^{\mathrm{BS}}_{k}[i]\delta_{\mathrm{t}}}{\varrho_{\mathrm{BS}}}\geq\sum\limits_{i=1}^{n-1}l^{\mathrm{BS}}_{k}[i],~{}k\in\mathcal{K},\forall n\in\mathcal{N}_{1},$ (19)
where $\varrho_{\mathrm{BS}}$ and $\varrho_{\mathrm{ua}}$ represent the number of CPU cycles needed for processing 1-bit of task-input data at the BS and UAV, respectively, and $\mathcal{N}_{1}\triangleq\{2,3,\cdots,N\}$ is a subset of $\mathcal{N}$. Therefore, the total amount of the completed task-input data for all users within the whole period can be calculated as
@@ III-A Optimization Problem Formulation
  • the resource allocation variables in $\mathbf{L}\triangleq\{l^{\mathrm{ua}}_{k}[n],l^{\mathrm{BS}}_{k}[n]$, $f^{\mathrm{ua}}_{k}[n],f^{\mathrm{BS}}_{k}[n],k\in\mathcal{K},n\in\mathcal{N}\}$;
  • the user scheduling variables for task offloading in $\boldsymbol{\zeta}\triangleq\{\zeta_{k}[n],k\in\mathcal{K},n\in\mathcal{N}\}$;
  • the passive beamforming variables of STAR-RIS in $\boldsymbol{\Upsilon}\triangleq\{\boldsymbol{\Theta}_{\mathrm{r}}[n]$, $\boldsymbol{\Theta}_{\mathrm{t}}[n],n\in\mathcal{N}\}$;
  • the UAV trajectory variables in $\mathbf{Q}\triangleq\{\mathbf{q}_{\mathrm{ua}}[n],n\in\mathcal{N}\}$.
@@
$\quad(\beta_{\mathrm{r}}^{m}[n])^{2}+(\beta_{\mathrm{t}}^{m}[n])^{2}=1,~{}\forall m\in\mathcal{M},n\in\mathcal{N},$ (26f)
$\quad\beta_{\mathrm{r}}^{m}[n],\beta_{\mathrm{t}}^{m}[n]\in(0,1),~{}\forall m\in\mathcal{M},n\in\mathcal{N},$ (26g)
$\quad\phi_{\mathrm{r}}^{m}[n],\phi_{\mathrm{t}}^{m}[n]\in[0,2\pi),~{}\forall m\in\mathcal{M},n\in\mathcal{N}.$ (26h)
Actually, the problem (III-A) is a non-convex problem due to the non-convexity of the objective function, and constraints (14), (15) and (26f), which is difficult to solve directly. To overcome this challenge, the alternative strategy is employed to divide the original optimization problem into three subproblems. The algorithm is designed by alternatively optimizing three variable subsets, which are respectively denoted as $\boldsymbol{\Xi}_{1}=\{\mathbf{L},\boldsymbol{\zeta}\}$, $\boldsymbol{\Xi}_{2}=\{\mathbf{L},\boldsymbol{\Upsilon}\}$ and $\boldsymbol{\Xi}_{3}=\{\mathbf{L},\mathbf{Q}\}$. More details of the algorithm design is given in the next subsection.
@@ III-B1 Designing $\boldsymbol{\Xi}_{1}$ with the given $\mathbf{Q}$ and $\boldsymbol{\Upsilon}$
First, we jointly optimize the resource allocation variable $\mathbf{L}$ and user scheduling variable $\boldsymbol{\zeta}$ with the given passive beamforming and UAV trajectory. In this case, the original problem (III-A) can be simplified as
$\max_{\boldsymbol{\Xi}_{1}}~{}\frac{L_{\mathrm{tol}}(\boldsymbol{\Xi}_{1})}{E_{\mathrm{tol}}(\boldsymbol{\Xi}_{1})}$
@@
Note that the optimization problem (III-B1) is still a non-convex problem because of the fractional objective function and the binary variable $\boldsymbol{\zeta}$. To tackle this problem, we first leverage the Dinkelbach’s algorithm [27, Chapter 3.2.1] to transform the fractional objective function . Specifically, the objective function in the $(l+1)$-th iteration of the Dinkelbach’s algorithm can be expressed as
$\widehat{\gamma}^{(l+1)}=L_{\mathrm{tol}}(\boldsymbol{\Xi}_{1})-\psi^{(l)}E_{\mathrm{tol}}(\boldsymbol{\Xi}_{1}),$ (28)
where $\psi^{(l)}$ can be updated with $\psi^{(l)}=\frac{L_{\mathrm{tol}}(\boldsymbol{\Xi}_{1}^{(l)})}{E_{\mathrm{tol}}(\boldsymbol{\Xi}_{1}^{(l)})}$.
In terms of the non-convex binary constraint (26d), we equivalently transform it into the following constraints:
$\sum_{k=1}^{K}\zeta_{k}[n]=1,0\leq\zeta_{k}[n]\leq 1,~{}\forall n\in\mathcal{N},k\in\mathcal{K},$ (29)
$\eta_{k}[n]=\zeta_{k}[n]-\zeta_{k}^{2}[n]=0,~{}\forall n\in\mathcal{N},k\in\mathcal{K}.$ (30)
@@
  • $\eta_{k}[n]\leq\widehat{\eta}_{k}[n]\big{(}\zeta_{k}[n],\zeta_{k}^{(t)}[n]\big{)}=\zeta_{k}[n]-\big{(}(\zeta_{k}^{(t)}[n])^{2}+2\zeta_{k}^{(t)}[n](\zeta_{k}[n]-\zeta_{k}^{(t)}[n])\big{)}$.
Thus, in the $(t+1)$-th inner loop iteration of the $(l+1)$-th iteration of the Dinkelbach’s algorithm, the optimization problem (III-B1) can be re-expressed
$\max_{\boldsymbol{\Xi}_{1}}~{}\widehat{\gamma}^{(l+1)}-\hat{\rho}\sum\limits_{n=1}^{N}\sum\limits_{k=1}^{K}\widehat{\eta}_{k}[n]\left(\zeta_{k}[n],\zeta_{k}^{(t)}[n]\right)$
@@
where $\hat{\rho}>0$ denotes the introduced penalty coefficient. The problem (III-B1) is a convex optimization problem and can be directly solved by the existing tool such as CVX.
To solve problem (III-B1), we propose a two-tier iterative algorithm which is summarized as Algorithm 1. The primary objective of the outer loop is to modify the fractional objective function using the Dinkelbach’s algorithm. On the other hand, the inner loop aims to ensure that the binary constraint is met by progressively increasing the penalty coefficient with $\hat{\rho}=\omega\hat{\rho}$, where $\omega>1$ is the scaling factor.

### v3

@@ II-B Task Offloading and Computation Model
where $p$ denotes the unified transmitted power of users, $B$ is the bandwidth of the system, $\sigma^{2}_{\mathrm{ua}}$ and $\sigma^{2}_{\mathrm{BS}}$ respectively denote the noise power at the BS and the UAV. In addition, $\boldsymbol{\Theta}_{\kappa}[n]=\operatorname{Diag}\big{\{}\beta^{1}_{\kappa}[n]e^{j\phi^{1}_{\kappa}[n]},\cdots,\beta^{m}_{\kappa}[n]e^{j\phi^{m}_{\kappa}[n]},\cdots,\beta^{M}_{\kappa}[n]e^{j\phi^{M}_{\kappa}[n]}\big{\}}$ is the matrix of the STAR-RIS’s coefficients with $\kappa\in\{\mathrm{r},\mathrm{t}\}$ representing the reflected or transmitted coefficients of the STAR-RIS, where the amplitudes $\beta^{m}_{\kappa}[n]$ and phases $\phi^{m}_{\kappa}[n]$ of the STAR-RIS should satisfy: $\beta^{m}_{\mathrm{r}}[n],~{}\beta^{m}_{\mathrm{t}}[n]\in(0,1]$ $(\beta^{m}_{\mathrm{r}}[n])^{2}+(\beta^{m}_{\mathrm{t}}[n])^{2}=1$, $\phi^{m}_{\mathrm{r}}[n],\phi^{m}_{\mathrm{t}}[n]\in[0,2\pi],~{}\forall m\in\mathcal{M}$.
It is assumed that the ground users are with limited resources for local computing and thus users’ tasks have to be offloaded to the BS and UAV for computing [^1] The local processing at the users is disregarded due to the following reasons: (i) For certain fundamental devices such as smart sensors, edge cameras, surveillance systems, and health monitoring devices, carrying out local computing tasks may pose various challenges. (ii) In situations where power resources are severely limited and not promptly replenished, devices with limited computing capabilities may opt to suspend local processing in order to conserve power. This is done with the aim of maximizing the device’s operational time and ensuring the continuous functioning of essential features.. Let $l^{\mathrm{BS}}_{k}[n]$ and $l^{\mathrm{ua}}_{k}[n]$ denote the number of the offloaded bits that need to be computed at the BS and UAV for user $k$ in the $n$-th time slot, respectively. Due to the fact that the BS and UAV can only deal with the tasks they have received, the following constraints should be satisfied
$\delta_{\mathrm{t}}\zeta_{k}[n]R^{\mathrm{ua}}_{k}[n]\geq l^{\mathrm{ua}}_{k}[n],~{}\forall k\in\mathcal{K},~{}n\in\mathcal{N},$ (14)
$\delta_{\mathrm{t}}\zeta_{k}[n]R^{\mathrm{BS}}_{k}[n]\geq l^{\mathrm{BS}}_{k}[n],~{}\forall k\in\mathcal{K},~{}n\in\mathcal{N}.$ (15)
In order to ensure that every user’s minimum computational need is met, we implement the following QoS constraints
@@
where $L_{k}$ is the minimum computing task requirement of the $k$-th user.
Let $f_{k}^{\mathrm{BS}}[n]$ and $f_{k}^{\mathrm{ua}}[n]$ respectively denote the computation frequency allocated by the BS and UAV for the $k$-th user at the $n$-th time slot. In addition, considering the processing causality, it is assumed that the BS and the UAV solely receive the offloaded tasks without carrying out any computations within the first time slot, and the users cease the act of offloading their tasks at the last time slot, i.e., $f^{\mathrm{BS}}_{k}[1]=f^{\mathrm{ua}}_{k}[1]=l^{\mathrm{BS}}_{k}[N]=l^{\mathrm{ua}}_{k}[N]=0$ for $k\in\mathcal{K}$. In order to guarantee that all users’ offloaded task-input data can be completely computed within the give mission time $T$, we have the following information-causality constraints:
@@
$\sum\limits_{i=2}^{n}\frac{f^{\mathrm{ua}}_{k}[i]\delta_{\mathrm{t}}}{\varrho_{\mathrm{ua}}}\geq\sum\limits_{i=1}^{n-1}l^{\mathrm{ua}}_{k}[i],~{}\forall k\in\mathcal{K},~{}n\in\mathcal{N}_{1},$ (18)
$\sum\limits_{i=2}^{n}\frac{f^{\mathrm{BS}}_{k}[i]\delta_{\mathrm{t}}}{\varrho_{\mathrm{BS}}}\geq\sum\limits_{i=1}^{n-1}l^{\mathrm{BS}}_{k}[i],~{}\forall k\in\mathcal{K},~{}n\in\mathcal{N}_{1},$ (19)
where $F_{\mathrm{BS}}$ and $F_{\mathrm{ua}}$ are the maximum CPU frequency at the BS and the UAV, respectively. $\varrho_{\mathrm{BS}}$ and $\varrho_{\mathrm{ua}}$ represent the number of CPU cycles needed for processing 1-bit of task-input data at the BS and UAV, respectively, and $\mathcal{N}_{1}\triangleq\{2,3,\cdots,N\}$ is a subset of $\mathcal{N}$. Therefore, the total amount of the completed task-input data for all users within the whole period can be calculated as
@@ III-A Optimization Problem Formulation
  • the resource allocation variables in $\mathbf{L}\triangleq\{l^{\mathrm{ua}}_{k}[n],l^{\mathrm{BS}}_{k}[n]$, $f^{\mathrm{ua}}_{k}[n],f^{\mathrm{BS}}_{k}[n],~{}k\in\mathcal{K},~{}n\in\mathcal{N}\}$;
  • the user scheduling variables for task offloading in $\boldsymbol{\zeta}\triangleq\{\zeta_{k}[n],~{}k\in\mathcal{K},~{}n\in\mathcal{N}\}$;
  • the passive beamforming variables of STAR-RIS in $\boldsymbol{\Upsilon}\triangleq\{\boldsymbol{\Theta}_{\mathrm{r}}[n]$, $\boldsymbol{\Theta}_{\mathrm{t}}[n],~{}n\in\mathcal{N}\}$;
  • the UAV trajectory variables in $\mathbf{Q}\triangleq\{\mathbf{q}_{\mathrm{ua}}[n],~{}n\in\mathcal{N}\}$.
@@
$\quad(\beta_{\mathrm{r}}^{m}[n])^{2}+(\beta_{\mathrm{t}}^{m}[n])^{2}=1,~{}\forall m\in\mathcal{M},~{}n\in\mathcal{N},$ (26f)
$\quad\beta_{\mathrm{r}}^{m}[n],\beta_{\mathrm{t}}^{m}[n]\in(0,1],~{}\forall m\in\mathcal{M},~{}n\in\mathcal{N},$ (26g)
$\quad\phi_{\mathrm{r}}^{m}[n],\phi_{\mathrm{t}}^{m}[n]\in[0,2\pi),~{}\forall m\in\mathcal{M},~{}n\in\mathcal{N}.$ (26h)
Actually, the problem (26) is a non-convex problem due to the non-convexity of the fractional objective function, and constraints (14), (15) and (26f), which is difficult to solve directly. To effectively handle the optimization problem with the fractional objective function, Dinkelbach’s algorithm, as one of the most popular fractional programming algorithm, will be leveraged. According to its core principle, we first transform the problem (26) as
$\max_{\boldsymbol{\Gamma},\psi}~{}~{}L_{\mathrm{tol}}-\psi E_{\mathrm{tol}},$
[…]
where $\psi$ denotes the introduced auxiliary variable, $\boldsymbol{\Gamma}=\{\mathbf{L},\boldsymbol{\zeta},\boldsymbol{\Upsilon},\mathbf{Q}\}$. If $\psi^{\mathrm{opt}}=\frac{L_{\mathrm{tol}}(\boldsymbol{\Gamma}^{\mathrm{opt}})}{E_{\mathrm{tol}}(\boldsymbol{\Gamma}^{\mathrm{opt}})}$ denotes the optimal objective value of the optimization problem (26), $L_{\mathrm{tol}}(\boldsymbol{\Gamma})-\psi^{\mathrm{opt}}E_{\mathrm{tol}}(\boldsymbol{\Gamma})\leq 0$ always holds for any $\boldsymbol{\Gamma}$ and the equality occurs exclusively when the optimal solution $\boldsymbol{\Gamma}^{\mathrm{opt}}$ of problem (26) is reached. Therefore, the optimization problems (27) and (26) will converge to the same optimal solution when $\psi=\psi^{\mathrm{opt}}$. This equivalence allows us to resolve the transformed problem (27) to attain the optimal solution for the original problem (26). However, it is challenging to obtain $\psi^{\mathrm{opt}}$ in advance. To address this challenge, the Dinkelbach’s algorithm opts to iteratively update the $\psi$ based on the solution of the transformed problem (27) to gradually achieve the optimal solution of the optimization problem (26). Hence, the Dinkelbach’s algorithm demonstrates considerable potential in efficiently tackling optimization problems with fractional objective functions. More details about the Dinkelbach’s algorithm are presented in [41, Chapter 3.2.1].
On the basis of the analysis above, the original problem (26) can be transformed as the following problem in the $(l+1)$-th iteration of the Dinkelbach’s algorithm, which is given by
$\max_{\boldsymbol{\Gamma}}~{}~{}L_{\mathrm{tol}}-\psi^{(l)}E_{\mathrm{tol}},$
[…]
where $\psi^{(l)}=\frac{L_{\mathrm{tol}}^{(l)}}{E_{\mathrm{tol}}^{(l)}}$. Note that the optimization problem (28) is still a non-convex optimization problem due to the significant coupling among variables in constraints (14) and (15), as well as the equality constraint (26f). To overcome this challenge, the alternative strategy is employed to divide the optimization problem (28) into three subproblems. The algorithm is designed by alternatively optimizing three variable subsets, which are respectively denoted as $\boldsymbol{\Xi}_{1}=\{\mathbf{L},\boldsymbol{\zeta}\}$, $\boldsymbol{\Xi}_{2}=\{\mathbf{L},\boldsymbol{\Upsilon}\}$ and $\boldsymbol{\Xi}_{3}=\{\mathbf{L},\mathbf{Q}\}$. More details of the algorithm design is given in the next subsection.
@@ III-B1 Designing $\boldsymbol{\Xi}_{1}$ with the given $\mathbf{Q}$ and $\boldsymbol{\Upsilon}$
First, we jointly optimize the resource allocation variable $\mathbf{L}$ and user scheduling variable $\boldsymbol{\zeta}$ with the given passive beamforming and UAV trajectory. In this case, the original problem (26) can be simplified as
$\max_{\boldsymbol{\Xi}_{1}}~{}L_{\mathrm{tol}}(\boldsymbol{\Xi}_{1})-\psi^{(l)}E_{\mathrm{tol}}(\boldsymbol{\Xi}_{1}),$
@@
It is important to highlight that the optimization problem (29) is a non-convex problem because of the binary variable $\boldsymbol{\zeta}$. To address this optimization problem, the non-convex binary constraint (26d) is equivalently transformed as the following constraints:
$\sum_{k=1}^{K}\zeta_{k}[n]=1,0\leq\zeta_{k}[n]\leq 1,~{}\forall n\in\mathcal{N},~{}k\in\mathcal{K},$ (30)
$\eta_{k}[n]=\zeta_{k}[n]-\zeta_{k}^{2}[n]=0,~{}\forall n\in\mathcal{N},~{}k\in\mathcal{K}.$ (31)
@@
$\eta_{k}[n]\leq$ $\widehat{\eta}_{k}[n]\big{(}\zeta_{k}[n],\zeta_{k}^{(t)}[n]\big{)}=\zeta_{k}[n]-\big{(}(\zeta_{k}^{(t)}[n])^{2}$
$+2\zeta_{k}^{(t)}[n](\zeta_{k}[n]-\zeta_{k}^{(t)}[n])\big{)}.$ (32)
Thus, in the $(t+1)$-th inner loop iteration of the $(l+1)$-th iteration of the Dinkelbach’s algorithm, the optimization problem (29) can be re-expressed
$\max_{\boldsymbol{\Xi}_{1}}~{}L_{\mathrm{tol}}(\boldsymbol{\Xi}_{1})-\psi^{(l)}E_{\mathrm{tol}}(\boldsymbol{\Xi}_{1})$
$\quad~{}~{}~{}-\hat{\rho}\sum\limits_{n=1}^{N}\sum\limits_{k=1}^{K}\widehat{\eta}_{k}[n]\left(\zeta_{k}[n],\zeta_{k}^{(t)}[n]\right),$
@@
where $\hat{\rho}>0$ denotes the introduced penalty coefficient. The problem (33) is a convex optimization problem and can be directly solved by the existing tool such as CVX.
Algorithm 1: The Proposed Iterative Algorithm for Solving the Sub-problem (29)
1: Initialize the feasible point $\big{(}\mathbf{L}^{(l,0)},\boldsymbol{\zeta}^{(l,0)}\big{)}$; Define the
 tolerance accuracy thresholds as $\widehat{\varepsilon}$; Set the iteration index
$t$ = 0; Initialize $\hat{\rho}^{(0)}$
2: While $\widehat{v}>\widehat{\varepsilon}$ or $t=0$ do
3:  Solve the optimization problem (33) with the given
    $\boldsymbol{\zeta}^{(l,t)}$ and update $\big{(}\mathbf{L}^{(l,t+1)},\boldsymbol{\zeta}^{(l,t+1)}\big{)}$ with the obtained
    solutions.
4:  Calculate $\widehat{v}=\max\limits_{k\in\mathcal{K},n\in\mathcal{N}}\eta_{k}[n]$ based on the acquired
    Let $t=t+1$.
5: end while
6: Update $\big{(}\mathbf{L}^{(l+1,0)},\boldsymbol{\zeta}^{(l+1,0)}\big{)}$ with $\big{(}\mathbf{L}^{(l,t)},\boldsymbol{\zeta}^{(l,t)}\big{)}$.
To solve problem (29), we propose an iterative algorithm which is summarized as Algorithm 1. The primary objective aims to ensure that the binary constraint is met by progressively increasing the penalty coefficient with $\hat{\rho}=\omega\hat{\rho}$, where $\omega>1$ is the scaling factor.